# The Yōkai Learning Environment: Tracking Beliefs Over Space and Time

### v1

@@ Abstract
Developing collaborative AI hinges on Theory of Mind (ToM) – the ability to reason about the beliefs of others to build and maintain common ground. Existing ToM benchmarks, however, are restricted to passive observer settings or lack an assessment of how agents establish and maintain common ground over time. To address these gaps, we introduce the Yōkai Learning Environment (YLE) – a multi-agent reinforcement learning (RL) environment based on the cooperative card game Yōkai. In the YLE, agents take turns peeking at hidden cards and moving them to form clusters based on colour. Success requires tracking evolving beliefs, remembering past observations, using hints as grounded communication, and maintaining common ground with teammates. Our evaluation yields two key findings: First, current RL agents struggle to solve the YLE, even when given access to perfect memory. Second, while belief modelling improves performance, agents are still unable to effectively generalise to unseen partners or form accurate beliefs over longer games, exposing a reliance on brittle conventions rather than robust belief tracking. We use the YLE to investigate research questions in belief modelling, memory, partner generalisation, and scaling to higher-order ToM.
@@ 1 Introduction
Effective collaboration among partners requires common ground – the shared knowledge, beliefs, and assumptions that enable them to interpret each other’s actions and respond appropriately Clark (1996). When common ground cannot be reliably maintained, collaboration often breaks down Klein et al. (2005). To establish common ground, humans use Theory of Mind (ToM) Premack and Woodruff (1978) to infer others’ mental states. Consequently, enabling computational agents to perform ToM reasoning has been identified as a critical step toward effective human-AI collaboration Klien et al. (2004); Dafoe et al. (2020, 2021).
Research in cognitive science suggests that human ToM integrates multiple reasoning processes – including memory, spatial navigation, and mentalising Suddendorf and Corballis (2007); Mitchell (2009). Prior work in computational ToM has focused on the latter, for example, by predicting others’ behaviour from an omniscient observer’s perspective Baker et al. (2011); Rabinowitz et al. (2018); Gandhi et al. (2021). However, humans are rarely passive observers of a scene where they solely predict behaviour Sclar et al. (2022). Instead, they are active participants and must adjust their ToM reasoning to the interactive setting in which their own actions change the environment, influence the behaviour of others, and vice versa. This interactive structure necessitates tracking how common ground – the mutually recognised shared experience built through interaction Tomasello (2010) – evolves during collaboration.
The goal of this paper is thus to
propose an environment in which effective collaboration between computational agents requires building common ground by tracking beliefs over space and time.
@@
Figure 1: The YLE poses a challenging ToM reasoning task for cooperative agents in a dynamic environment. Agents cannot observe all cards in a single game. Successful play requires agents to remember and reason about others and their own mental states. In the example, it’s Refer to caption ’s turn. Refer to caption observes the colour of two cards (1 and 2) privately. Refer to caption recalls from earlier that card 1 is blue. When Refer to caption moves card 1 next to card 2, Refer to caption can infer that the second card is likely blue as well. Refer to caption and Refer to caption can now assume this is common ground going forward.
@@
This is the Yōkai Learning Environment (YLE) [^1] Code available at https://git.hcics.simtech.uni-stuttgart.de/public-projects/Yokai_env. – a novel multi-agent reinforcement learning (RL) environment based on the collaborative card game Yōkai. In the YLE, players must collaborate to group face-down cards by colour over multiple rounds without direct communication (see Figure 1, left). Each round, one player privately observes the colour of two cards, moves one of them to a new position on the table, and may place a hint card to communicate information about unobserved card colours to others (see Figure 2). Importantly, players may end the game early for a higher reward, but at the risk of losing the game. This makes it a high-stakes, all-or-nothing decision that requires strong ToM reasoning to infer card colours from partners’ actions rather than first-hand observation.
The design of YLE creates a uniquely challenging coordination task. Agents must: (1) remember the colours of moving cards to avoid redundant checks, (2) infer the beliefs and knowledge of others to establish and maintain common ground, (3) interpret teammates’ actions and hints as implicit belief updates, (4) use hint cards as grounded communication, and (5) assess when to end early based on inferred shared knowledge. These challenges require belief tracking over both space (shifting card locations) and time (multi-round gameplay with limited observability), and together reflect key components of ToM Premack and Woodruff (1978); Clark (1996); Pillow (1989); Southgate et al. (2007); Tomasello (2010) in a dynamic, interactive setting. Effectively establishing common ground requires aligning on ad hoc conventions for grouping cards with partners. Prior environments, such as Hanabi Bard et al. (2020), were used to study zero-shot coordination and ToM, but lack the spatial-temporal structure, evolving belief dynamics, and the challenge of ending early based on inferred knowledge.
We evaluate a range of agents on YLE and highlight key limitations in their ability to track common ground in this novel spatial-temporal belief reasoning task. Our experiments show that even with access to perfect memory, agents struggle to collaborate effectively. Among agents without perfect memory, those with explicit memory modules perform best, and performance drops further in four-player settings. While belief representation learning improves performance, agents fail to generalise to new partners or maintain accurate beliefs over longer time horizons. Instead, they rely on brittle conventions rather than robust and generalisable belief tracking. Our contributions are:
@@
    A fast JAX-based YLE – a new cooperative multi-agent RL benchmark for belief tracking, common ground, and ToM, training at hundreds of thousands of steps per second.
@@
    An empirical analysis showing that agents fail to coordinate effectively in YLE, and in particular fail to end games early – a behaviour humans achieve reliably.
@@
    A demonstration that belief representation learning improves performance, but agents still fail to generalise to novel partners or maintain accurate beliefs over time.
  4. 4.
    An evaluation of YLE in a four-player setting, revealing the increased difficulty of maintaining common ground when high-order ToM is required.
2 Yōkai
@@
Figure 2: A sample round in nine-card YLE. Agent Refer to caption first observes two cards (one blue, one green), moves the blue card, and finishes by either (4a) revealing or (4b) placing a hint. Refer to caption places a hint. Then, Refer to caption can start its turn. Refer to caption chooses to end the game. Ending the game will return the final reward based on the outcome. This game is lost because while blue cards are clustered, green and red ones are not.
@@
Yōkai [^2] Detailed rules for Yōkai are available at http://boardgame.bg/yokai%20rules.pdf. is a turn-based cooperative game in which players must sort face-down cards into same-colour clusters to win. A key requirement is that all cards must be connected to at least one other card on their sides. Classically, the game is played by two to four players with 16 cards of four colours each. Complementing these 16 Yōkai cards, additional hint cards feature one or multiple colours and are used to hint at the colour of a face-down card underneath them. They are the only allowed communication mechanism in the game. The number of players determines the number of hint cards, which varies between seven and 10 (see Appendix B for further details). Hint cards are initially face-down and must be revealed before being placed face-up on a Yōkai card. Importantly, this Yōkai card is locked in place and can no longer be moved or observed. Since most hints can be multi-coloured, their interpretation is left to the other agents, depending on the context and history.
@@
A typical turn in Yōkai consists of four steps (see Figure 2). The current player first observes two cards privately, then moves one card, and then either (a) reveals or (b) places a face-up hint card. Consequently, the available legal moves depend on the specific card to be moved. Figure 11, while Figure 10 show legal moves and card configurations. The game ends when all hint cards have been revealed and placed. Optionally, players can end the game prematurely instead of observing cards at the beginning of their turn. The final game score is based on the final card configuration:
@@
$\begin{split}S=&\;5N_{\text{hints\_face\_down}}+2N_{\text{hints\_not\_placed}}+N_{\text{hints\_correct}}\\ &-N_{\text{hints\_wrong}}.\end{split}$ (1)
@@
Players are encouraged to end the game early to achieve a higher score, since unused hint cards contribute most to the reward. However, the earlier a player ends a game, the less information they can acquire first-hand and thus must infer knowledge from others’ behaviour (see Figure 12 for an example). As such, the final game score measures players’ joint ToM reasoning abilities. In particular, the rate at which players successfully end the game early reflects how much they rely on belief inference, making it a powerful new implicit measure of ToM reasoning under uncertainty.
@@
Yōkai poses three key challenges to cooperative agents: First, achieving a high score requires agents to end a game as early as possible, resulting in a higher risk of failure. Second, ending the game early requires interpreting others’ actions and tracking their beliefs to form correct own beliefs over the colours of unobserved cards (see Figure 1). Third, agents must use hints efficiently as grounded communication, but a single misplaced hint can render the game unwinnable.
@@ 3 Preliminaries
We model Yōkai as a decentralised partially observable Markov decision process (Dec-POMDP) Oliehoek and Amato (2016), formally described as a tuple $\mathcal{M}=\langle\mathcal{N},\mathcal{S},\Omega,O,A,\mathcal{R},\mathcal{T},\gamma\rangle$. $\mathcal{N}$ is the set of agents where $i\in\mathcal{N}$, $1\leq i\leq|\mathcal{N}|$ denotes one specific agent and i-i denotes all other agents. Set $\mathcal{S}$ denotes the possible true states of the environment, with $s\in\mathcal{S}$ being a single state. At timestep $t$ an agent in state $s_{t}$ receives a partial observation $o^{i}_{t}\in\Omega$ through the observation function $O(s_{t},i)$. The action-observation history per agent is denoted by $\tau^{i}$. Following Sclar et al. Sclar et al. (2022), an agent with no ToM chooses an action $a^{i}\in A$ only based on their observation of the environment $\pi^{i}(a^{i}|o^{i}_{t})$. Agents with zeroth-order ToM incorporate memory, conditioning their policy on a hidden state $h^{i}_{t}$ in addition to their current observation $\pi^{i}(a^{i}|o^{i}_{t},h^{i}_{t})$. Furthermore, agents that reason about the mental states of others estimate their hidden states $\hat{h}_{t}^{(j)}$ for $j\neq i$ and condition their action on all $\hat{h}_{t}^{(-i)}$: $\pi^{i}(a^{i}|o^{i}_{t},h^{i}_{t},\hat{h}_{t}^{-i})$. Based on the joint action performed by all agents $\bm{a}=(a_{1},\dots,a_{n})$ the environment returns a shared reward $r_{t}=R(s_{t},\bm{a})$ and then produces a new state based on the state transition function $\mathcal{T}$. Finally, using the discount factor $\gamma^{t}$, the agents optimise the discounted return of a trajectory given by $R(\tau)=\sum_{t}\gamma^{t}r_{t}$.
@@ 4 The Yōkai Learning Environment
The Yōkai Learning Environment implements the game logic with two exceptions: First, we support configurable game sizes (e.g., a 3x3 version with nine cards and three colours). We refer to it as 3x3 while the original game is referred to as 4x4. The 3x3 two-player game has four hint cards: one one-colour hint and three two-colour hints. Second, our game is restricted to a $g\times g$ grid to efficiently compute legal moves. We pick $g$ large enough not to limit gameplay in practice and confirm this experimentally, see Appendix B. Our implementation is based on JaxMARL Rutherford et al. (2023) and is fully GPU-accelerated. It allows end-to-end learning on a single GPU at hundreds of thousands of steps-per-second, see Table 3.
@@
We opted to model Yōkai as a graph since it is a natural formulation of the relational nature of the environment and allows for an efficient implementation in JAX. At time $t$, the state is a graph $G_{t}(V,A)$ with cards and hints as vertices $V=Y\cup H$ and adjacency matrix $A$. Within $G_{t}$, cards are vertices $V=Y\cup H$ and connected to neighbours via edges represented by an adjacency matrix $A$. Cards and hints are identified by their index $0\leq i<|Y|$ or $0\leq j<|H|$. Hint cards are not connected to any card. For example, in the 3x3 version with nine cards and four hints, colours (e.g., blue, green, red) and hints are randomly assigned each episode, and the graph $A\in\{0,1\}^{13\times 13}$ is updated when cards move. With this setup, moving a card $i$ boils down to obtaining $A^{\prime}$ from $A$ via $A^{\prime}=\text{move}(A,i,n)$ where in move the row and column vector at index $i$ of $A$ are replaced with a binary encoding of the new neighbours $n$. Additionally, a configuration is legal if all cards are connected, computed via a reachability matrix over $A$. Determining whether the game has been won then boils down to computing
$\bigwedge\limits_{c\in C}\texttt{isConnected}(A_{c})$ (2)
where $A_{c}$ denotes the submatrix of $A$ obtained by deleting all column and row vectors not associated with colour $c$, i.e. by finding the set of indices $\{j|0\leq j<|Y|\land Y[j]\neq c\}$ where $Y[j]$ denotes accessing the tuple $Y$ at index $j$. To compute the state-dependent set of legal move actions $A(s)\subseteq A$, we run $\texttt{isConnected}(A^{\prime})$ on all $A^{\prime}$ obtainable by moving a card to a position in the $g\times g$ field.
@@
Agents receive graph-based or image-like observations encoding card properties (e.g., colour, position, lock state) and hint features. Observations include both public (topology and hints) and private (card features) information, leading to asymmetric knowledge. Graph observations encode cards as an adjacency matrix $A_{t}\in\{0,1\}^{|Y|\times|Y|}$, and their properties as node features $F_{t}\in\mathbb{R}^{|Y|+|H|\times f}$, where $f$ is the feature dimension. In image-like observations, card features $F$ are embedded into a $g\times(g+1)\times f$ tensor based on their positions, while hint card features are embedded into column $g+1$. Agents observe which cards others inspect, but not their content. Partial observability forces agents to seek information and track shared knowledge.
@@
Actions in the YLE are represented by a categorical action space. Additional players increase the number of hints and, thereby, the number of possible actions. The action space includes observing pairs of cards, moving cards, placing or revealing hints, ending the game, and a no-op. The action space size is the sum of the number of cards to look at $|Y|$, the number of target positions any card can be moved to $N_{moves}=g^{2}|Y|$, and the number of hints $|H|$. The number of cards a hint can be placed onto $N_{hint\_moves}=|H|\cdot|Y|$, $1$ action for no operation and $1$ action to end the game. The exact size of the action space depends on the setup but ranges from $\sim 1,000$ to $\sim 3,000$ in practice, all details are in Appendix B.

### v2

@@ Abstract
The ability to cooperate with unknown partners is a central challenge in cooperative AI and widely studied in the form of zero-shot coordination (ZSC), which evaluates an algorithm by measuring the performance of independently trained agents when paired. The Hanabi Learning Environment (HLE) has become the dominant benchmark for ZSC, but recent work has achieved near-perfect inter-seed cross-play performance, limiting its ability to track algorithmic progress. We introduce the Yōkai Learning Environment (YLE) – an open-source multi-agent RL benchmark in which effective collaboration requires building common ground by tracking and updating beliefs over moving cards, reasoning under ambiguous hints, and deciding when to terminate the game based on inferred shared knowledge – features absent in the HLE, where beliefs are tied to hand slots and hints are truthful by rule. We evaluate the leading ZSC methods, including High-Entropy IPPO, Other-Play, and Off-Belief Learning, which achieve near-perfect inter-seed cross-play in the HLE, and show that in the YLE they exhibit persistent SP–XP gaps, degraded early-ending calibration, and weaker belief representations in cross-play, indicating failure to maintain consistent internal models with unseen partners. Methods that perform best in the HLE do not perform best in the YLE, indicating that progress measured on a single benchmark may not generalise. Together, these results establish YLE as a challenging new ZSC benchmark.
@@ 1 Introduction
Effective collaboration among partners requires common ground – the shared knowledge, beliefs, and assumptions that enable them to interpret each other’s actions and respond appropriately (Clark, 1996). Maintaining common ground requires reasoning about what a partner knows, believes, and intends – commonly referred to as Theory of Mind (ToM) (Premack & Woodruff, 1978). Improving ToM reasoning in AI agents has emerged as an important requirement toward effective human-AI collaboration (Klien et al., 2004; Dafoe et al., 2020; 2021). A necessary precondition for human-AI collaboration is zero-shot coordination (ZSC) (Hu et al., 2020) which focuses on designing algorithms that, when run independently multiple times, produce compatible policies.
The Hanabi Learning Environment (HLE) (Bard et al., 2020) is arguably the most widely used benchmark for ZSC and ToM (Hu et al., 2020; 2021; Fuchs et al., 2021; Cui et al., 2021; 2022; Muglich et al., 2022; 2025; Lauffer et al., 2025). However, close to perfect inter-seed cross-play performance has recently been reached (Sudhakar et al., 2025; Forkel et al., 2026), limiting its utility as a stress test for further ZSC progress. Consequently, the field would benefit from new benchmarks that can be used to track algorithmic advances without risking overfitting to the HLE.
@@
Figure 1: The YLE poses a challenging ToM reasoning task for ZSC. Agents cannot observe all cards in a single game. Successful play requires agents to reason about other agents’ knowledge and beliefs. In the example, it’s Refer to caption ’s turn. Refer to caption observes the colour of two cards (1 and 2) privately. Refer to caption recalls from earlier that card 1 is blue. When Refer to caption moves card 1 next to card 2, Refer to caption can infer that the second card is blue as well. Refer to caption and Refer to caption can use this common ground going forward.
@@
Our solution is the Yōkai Learning Environment (YLE)– a novel multi-agent reinforcement learning (RL) benchmark for ZSC based on the collaborative card game Yōkai. In the YLE, players must collaborate to group face-down cards by colour over multiple rounds without direct communication (see Figure 1, left). Each round, one player privately observes the colour of two cards, moves one of them to a new position, and may place a hint card to communicate information about card colours to others (see Figure 2). Hints in YLE can be multi-coloured and thus can be ambiguous. Furthermore hints can be placed on any of the yet unhinted board cards, including ones whose colour is not on the hint card. Hints are thus not required to be truthful.
Crucially, players may choose to end the game early for a substantially higher return than completing the game conservatively, but at the risk of receiving a return of zero. Because most of the achievable return lies in successfully ending as early as possible, this creates a high-stakes trade-off: players must commit early to maximise reward, yet doing so requires sufficiently accurate and well-calibrated beliefs about the evolving board state. Premature ending leads to catastrophic failure, while excessive caution yields suboptimal return.
To succeed in Yōkai, players must maintain accurate beliefs about dynamically moving cards, align interpretations of ambiguous hints and actions, and determine when evidence justifies early ending.
The design of YLE therefore creates a coordination problem under sparse, ambiguously grounded information and high-stakes commitment. Unlike Hanabi, where agents observe their partners’ hands through a small number of fixed card slots and hints are truthful by rule, YLE reveals information locally in space: no player ever observes the full board, hints may be multi-coloured and need not be truthful, and cards move across the grid. As a result, beliefs must be repeatedly rebound to dynamically moving cards while trying to end early under uncertainty about the true board configuration.
We show that ZSC algorithms, which achieve near-perfect inter-seed cross-play in HLE, exhibit persistent gaps between self-play (SP) and cross-play (XP) when evaluated in YLE, indicating coordination failures. We thus posit that (1) YLE is a timely new benchmark, as approaches that saturate HLE do not saturate YLE, and (2) YLE highlights limitations of existing ZSC methods that have so far been validated primarily in Hanabi. Our specific contributions are:
@@
    We introduce the Yōkai Learning Environment (YLE) as an open-source JAX-based benchmark for zero-shot coordination in which agents must build and maintain common ground through spatio-temporal belief tracking, ambiguous communication, and high-stakes early termination.
@@
    We provide comprehensive empirical evaluations of leading ZSC algorithms that achieve near-perfect inter-seed cross-play in Hanabi – including Other-Play (Hu et al., 2020), Off-Belief Learning (OBL) (Hu et al., 2021), and High-Entropy IPPO (Forkel et al., 2026). In the YLE, these methods exhibit persistent SP–XP gaps, degraded early-ending calibration, and reduced coordination. To our knowledge, this is the first implementation of OBL outside Hanabi.
@@
    We demonstrate that algorithm rankings change between HLE and YLE, suggesting that apparent progress in ZSC may be benchmark-specific.
2 Background: Yōkai
@@
Figure 2: A sample round in nine-card YLE. Refer to caption first observes two cards privately (one blue, one green), moves the blue card, and finishes by either (4a) revealing or (4b) placing a hint. Refer to caption places a hint. Then, Refer to caption can start their turn. Refer to caption chooses to end the game. Ending the game will return the final reward based on the outcome. This game is lost because green and red cards are not clustered.
Figure 3: Legal moves for the card with the blue dot. Cards can only be moved so that all cards remain connected via their sides. Cards with an orange dot cannot currently be moved.
@@
Yōkai [^1] Detailed rules for Yōkai are available at http://boardgame.bg/yokai%20rules.pdf. developed by Julien Griffon, is a turn-based cooperative game in which players must sort face-down cards into same-colour clusters to win. A key requirement is that all cards must be connected to at least one other card on their sides. Classically, the game is played by two to four players with 16 cards of four colours each. Complementing these Yōkai cards, additional hint cards feature one or multiple colours and are used to hint at the colour of a face-down card underneath them. Hint cards are the only allowed communication mechanism in the game, and their number depends on the number of players (see supplementary material; typically 4 – 10). These cards are initially face-down and must be revealed before being placed face-up on a Yōkai card. Importantly, a hinted Yōkai card can no longer be moved or observed. Since most hints can be multi-coloured, their interpretation is left to the other agents, depending on the context and history.
@@
A typical turn in Yōkai consists of four steps (see Figure 2). The current player first observes two cards privately, then moves one card, and then either (a) reveals or (b) places a face-up hint card. The available legal moves depend on the specific card to be moved (an example is shown in Figure 3). The game ends when all hint cards have been revealed and placed. Optionally, players can end the game early instead of observing cards at the beginning of their turn. The final game score is based on the final card configuration:
@@
$\begin{split}S=5N_{\text{hints\_face\_down}}+2N_{\text{hints\_not\_placed}}+N_{\text{hints\_correct}}-N_{\text{hints\_wrong}}.\end{split}$ (1)
@@
Players are encouraged to end the game early to achieve a higher score, since unused hint cards contribute most to the final score. However, the earlier a player ends a game, the less information they can acquire first-hand and instead from others’ behaviour (see Figure 4 for an example). As such, the final game score reflects players’ joint ToM reasoning abilities. In particular, the rate at which players successfully end the game early reflects how much they rely on belief inference, making it a powerful new implicit measure of ToM reasoning under uncertainty.
Figure 4: A second-order ToM reasoning example over multiple timesteps in 2-player YLE. Refer to caption beliefs that Refer to caption beliefs that they (Refer to caption ) knew that cards 1 and 2 are of the same colour. Even though Refer to caption never saw card 1 and Refer to caption never observed card 3, both now know where all blue cards are, that they are grouped and that both share this knowledge as part of their common ground. In the future, they now can potentially finish early as they have one less card that needs to be observed.
@@
Yōkai poses three key challenges to cooperative agents: First, achieving a high score requires agents to end a game as early as possible, resulting in a higher risk of failure. Second, ending the game early requires interpreting others’ actions and tracking their beliefs to form correct own beliefs over the colours of unobserved cards. Third, agents must use hints efficiently as grounded communication, but a single misplaced hint can render the game unwinnable.
@@ 3 Preliminaries
Definition 3.1 (Dec-POMDP (Oliehoek & Amato, 2016)).
[…]
$J_{\text{SP}}(\pi)=\mathbb{E}_{\tau_{T}\sim\pi}\left[\sum_{t=0}^{T-1}\gamma^{t}\mathcal{R}(s_{t},\bm{a}_{t})\right].$ (2)
Zero-Shot Coordination (ZSC) (Hu et al., 2020) aims to design algorithms that produce compatible policies in Dec-POMDPs, in that independently trained agents can coordinate. The compatibility between different joint policies $\pi_{1},...,\pi_{n}$ in an $n$-player Dec-POMDP is measured by the cross-play (XP) return, defined as in Forkel et al. (2026) as
$J_{\text{XP}}(\pi_{1},...,\pi_{n}):=\mathbb{E}_{\phi\sim\text{Perm}(|\mathcal{N}|)}\left[J_{\text{SP}}((\pi_{\phi(1)}^{1},...,\pi_{\phi(n)}^{n}))\right],$ (3)
where $\mathrm{Perm}(n)$ denotes the set of permutations of $\{1,\dots,n\}$. A common failure mode in ZSC is that agents learn partner-specific conventions that do not generalise to new partners, which results in XP returns being significantly lower than SP returns. Approaches to mitigate this SP-XP gap include Other-Play (OP) (Hu et al., 2020) and Off-Belief Learning (OBL) (Hu et al., 2021). Here, we focus on the commonly used inter-seed XP setting, where the implementation is fixed but random seeds differ (Hu et al., 2020; 2021; Cui et al., 2022; Muglich et al., 2025; Forkel et al., 2026).
@@ 4 The Yōkai Learning Environment
The Yōkai Learning Environment (YLE) implements a parametrised version of the collaborative card game Yōkai. We support configurable game sizes (e.g., 9C with nine cards and three colours, and 16C, the original game) and restrict play to a $g\times g$ grid to enable efficient computation of legal moves. The grid size $g$ is chosen large enough not to constrain gameplay in practice (see supplementary). The environment is implemented in JAX in JaxMARL (Rutherford et al., 2023) and supports end-to-end GPU training at hundreds of thousands of steps per second (see Table 1).
@@
We model Yōkai as a graph to reflect its relational structure and enable a efficient JAX implementation. At time $t$, the state is a graph $G_{t}(V,A)$ where vertices $V=Y\cup H$ represent cards and hints, and $A$ is the adjacency matrix encoding spatial neighbourhoods between cards. Hint cards are not connected to any vertex. Card colours and hints are randomly assigned at the beginning of each episode. Moving a card updates the corresponding row and column of $A$ to reflect its new spatial neighbours. A configuration is legal if all cards remain connected, which we verify via reachability on $A$. The game is won if, for each colour $c\in C$, all cards of colour $c$ form a connected component. Legal move actions are computed by evaluating connectivity after candidate moves.
@@
Agents receive either graph-based or image-like observations encoding card properties (e.g., colour, position, lock state, id) and hint features. Observations contain both public information (topology and hints) and private information (card features), leading to asymmetric knowledge. In the graph representation, cards are encoded by an adjacency matrix $A_{t}\in\{0,1\}^{|Y|\times|Y|}$ and node features $F_{t}\in\mathbb{R}^{(|Y|+|H|)\times f}$. In the image representation, card features are embedded into a $g\times(g+1)\times f$ image-like tensor according to spatial position. Agents observe which cards others inspect, but not their content, requiring belief tracking over time.
@@
Actions are represented categorically and include observing cards, moving cards, placing or revealing hints, ending the game, and a no-op. The action space size depends on the configuration and scales with the number of cards $|Y|$, grid positions $g^{2}$, and hints $|H|$. In practice, it ranges from approximately 1,0001{,}000 to 3,0003{,}000 actions (see supplementary). Invalid actions are masked at each timestep. Even with masking, this action space is substantially larger than those of widely used cooperative MARL and ZSC benchmarks (Bard et al., 2020; Carroll et al., 2019; Gessler et al., 2025) and therefore challenges existing algorithms with a larger exploration problem.